# Proximal Gradient Dynamics: Monotonicity, Exponential Convergence, and Applications

### v1

@@ IV Polyak-Łojaciewicz conditions for proximal gradient dynamics
    Note that if $V(x(0))$ is infinity, there is nothing to show. If $V(x(0))$ is finite, it stays finite for all future timesteps. Similar to the proof of Theorem 4, consider the Dini derivative of ${V(x(t))=f(x(t))+g(x(t))-F^{*}}$. Using the arguments presented in Theorem 4 upto Equation (11),
    $D^{+}V(x)\leq\nabla f(x)^{\top}(z-x)+s_{x}^{\top}(-x+z).$
    where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$, and $s_{x}\in\partial g(x)$. From the convexity of $g$, we have,
    $D^{+}V(x)\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$
    Adding and subtracting $\frac{1}{2\alpha}\|x-z\|_{2}^{2}$,
    $D^{+}V(x)$ $\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x)$
    $\quad+\frac{1}{2\alpha}\|x-z\|_{2}^{2}-\frac{1}{2\alpha}\|x-z\|_{2}^{2}.$ (13)
    Note that
[…]
    $\quad+\langle\nabla f(x),u-x\rangle\bigr{\}}$
[…]
    Therefore,
    $\nabla f(x)^{\top}(z-x)+g(z)-g(x)+\frac{1}{2\alpha}\|x-z\|_{2}^{2}$
    $\quad=\min_{y}\left(\nabla f(x)^{\top}(y-x)+g(y)-g(x)+\frac{1}{2\alpha}\|x-y\|_{2}^{2}\right)$
    $\quad=-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha)$
    Using this equality in (Proof:),
    $D^{+}V(x)$ $\leq-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha)-\frac{1}{2\alpha}\|x-z\|_{2}^{2}$ (14)
    Using the proximal PL condition (7),
    $D^{+}V(x)$ $\leq-\alpha\mu(F(x)-F(x^{\star}))-\frac{1}{2\alpha}\|x-z\|_{2}^{2}$ (15)
    Now, the second term may be bounded above by 00. Using Grönwall’s lemma we get our required result. ∎
@@
The PL condition in continuous time is only of value when $V(x(0))$ is finite. In discrete time, the proximal gradient algorithm forces a finite value for $V(\cdot)$ in one timestep. In general, in continuous time, one cannot characterize the time taken for $V(x(\cdot))$ to become finite.
Remark 13:
In [12, Theorem 3], the forward-backward envelope is shown to be globally exponentially stable under the PL conditions. The forward-backward envelope is an upper bound to $f(z)+g(z)$, under the assumption of $L-$smoothness of $f$, where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$. In contrast, here we are interested in the value of $f(x)+g(x)$. Identifying an upper bound on $f(z)+g(z)$ is particularly useful when $g(\cdot)$ takes nonfinite values but also does not reflect whether the current state is feasible.
@@ IV-A Extensions to time-varying optimization
In recent literature [8], variations of the classic problem (3) with time-varying parameters have also been considered. In this case, the goal is to define dynamics that track the minimizer of the optimization problem,
@@
Under Lipschitz constraints on $\theta$, one may define bounds for the difference of the current cost and minimum cost, when the PL condition is true.
@@
    Following the arguments of Theorem 11, the first term is bounded above by $-\alpha\mu(V(x,t))$. Since $F$ is Lipschitz in $\theta$, $\|\nabla_{\theta}F(x)\|\leq\ell_{\theta}$. Through an application of Cauchy Schwarz and triangle inequality, the second term may be bounded by $2\ell_{\theta}\|\dot{\theta}\|$ The final term is 00 due to the optimality of $x^{\star}$. In summary, we have,
@@ V Applications and Examples
In this section, we discuss a few applications of the proximal gradient dynamics.
V-A Least Squares with $\ell_{1}$ regularization
Through the proximal gradient dynamics, one may solve the $\ell_{1}$ regularized least squares problem, otherwise known as the Least Absolute Shrinkage and Selection Operator (LASSO). The problem is given by,
@@ V-A Least Squares with $\ell_{1}$ regularization
where $A\in\mathbb{R}^{m\times n},\;u\in\mathbb{R}^{m}$ The corresponding proximal gradient dynamics, referred to as the “firing rate competitive network” [6] are given by,
@@
    The LASSO problem is known to satisfy the PL condition [14]. The result follows from Theorems 11 and 14. ∎
V-B Quadratic optimization with polytopic constraints
Next, we consider another classic problem that is of interest in several domains. Consider the optimization problem,
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{\top}Qx-u^{\top}x\quad\text{s.t.}\quad Ax\leq b,\;Dx=c.$ (20)
where $Q\in\mathbb{R}^{n\times n}$ is positive definite, and other matrices are appropriately sized. Specific versions of this problem have been implemented both as neural circuits [4] and using analog circuits [20]. Further, it also represents the Fenchel dual to the cost function of support vector machines (SVMs) [14]. The corresponding proximal gradient dynamics are given by,
$\dot{x}=-x+\mathrm{proj}_{C}(x-\alpha Qx+\alpha u),$ (21)
where $C=\{x\;|\;Ax\leq b,\;Dx=c\}$.
Corollary 16:
Under the proximal gradient dynamics (21) for the constrained quadratic optimization problem (20), if the initial point is within the feasible set, and $\alpha=1/\lambda_{\min}(Q)$, the cost function decays to the optimal value exponentially with rate 1.
  •       Proof:
    Since we begin in the feasible set, the initial cost is finite. Therefore, by Theorem 9, the cost function converges. Since $f(x)$ is strongly convex, and the constraints set is convex, the proximal PL condition is true [14], with $\alpha=1/\lambda_{\min}(Q)$, and $\mu=\lambda_{\min}(Q)$. By applying Theorem 11, the rate of convergence for the cost function is $1$. ∎
V-C Matrix based problems
Finally, we consider matrix optimization problems. In this setting, the analogue of sparse vectors are low rank matrices. However, similar to the $\ell_{0}$ norm, imposing a low rank constraint leads to nonconvexity. A known convex relaxation for the $\ell_{0}$ norm is the trace norm. The proximal operator for $g(X)=\|X\|_{*}$ is
@@ V-C Matrix based problems
$\dot{X}=-X+\operatorname{S}_{\alpha\lambda}\bigl{(}{X-\alpha\sum_{i=1}^{k}(\operatorname{tr}(A_{i}^{\top}X)-y_{i})A_{i}}\bigr{)},$ (24)
A second popular matrix based optimization problem is the matrix factorization problem. This problem and its generalizations appear in overparameterized linear neural networks with mean square error loss [21]. In general, this problem is nonconvex. To promote a low-rank solution, it is customary to add terms proportional to the nuclear norm and write the problem as:
@@ VI Discussion
In this letter, we discuss some convergence properties of the proximal gradient dynamics. We prove in Theorem 4 that the cost function of the associated optimization problem is nonincreasing for finite-valued regularizers. For non-finite-valued regularizers, we are able to show that the cost is nonincreasing in Theorem 9. Next, we extend the PL conditions to continuous-time proximal gradient dynamics, explicitly showing that the cost function decays exponentially under the PL condition. We use this result to consider the case of parametric optimization problems with time-varying parameters, and provide a cost function tracking bound.
An interesting direction of future work is the application of proximal gradient dynamics to design continuous-time solvers, such as analog circuits and biologically plausible neural networks. Here, we discuss constrained optimization, LASSO problem and matrix recovery and factorization. However, these dynamics may also allow a generalized nonsmooth analysis of systems based on energy landscapes. A system that is currently studied under this framework is the human brain [11]. While previous work [19] suggests that the algorithms we design for modeling a brain must be gradient-based, due to these dynamics, one could design biologically plausible algorithms for a nonsmooth energy landscape.

### v2

@@ IV Exponential convergence of the cost
    The proof is deferred to Appendix -B
Corollary 10:
If the proximal PL inequality is true for a certain value of $\alpha$, it will also be true for all $\alpha^{\prime}\in(0,\alpha]$. This follows from the fact that $\mathcal{D}_{g}(x,\alpha^{\prime})\geq\mathcal{D}_{g}(x,\alpha)$.
@@
In [16, Theorem 3], the forward-backward envelope is shown to be globally exponentially stable under the PL conditions. The forward-backward envelope is an upper bound to $f(z)+g(z)$, under the assumption of $L-$smoothness of $f$, where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$. In contrast, here we are interested in the cost function. The cost function in many optimization problems has some physical interpretation, and hence this result is more intuitive.
@@ IV-A Extensions to time-varying optimization
In recent literature [12, 1, 18], variations of the classic problem (3) with time-varying parameters have also been considered. In typical applications, time-varying parameters arise from the evolution of the environment. In continuous time, the goal is to define dynamics that track the minimizer of the optimization problem,
@@
This formulation is generalizes the typical online optimization setup where $\theta(t)=t$. Under Lipschitz constraints on $\theta$, one may define bounds for the difference of the current cost and optimal cost. The difference of the cost incurred and the optimal cost is termed regret, and is a means to evaluate the performance of algorithms. Due to our exponential bound on the cost we may bound the regret at each time.
@@
    Following the arguments of Theorem 9, the first term is bounded above by $-\alpha\mu(V(x,t))$. Since $F$ is Lipschitz in $\theta$, $\|\nabla_{\theta}F(x)\|\leq\ell_{\theta}$. Through an application of Cauchy Schwarz and triangle inequality, the second term may be bounded by $2\ell_{\theta}\|\dot{\theta}\|$ The final term is 00 due to the optimality of $x^{\star}$. In summary, we have
@@ V Applications and Examples
Example #1: Least Squares with $\ell_{1}$ regularization
Through the proximal gradient dynamics, one may solve the $\ell_{1}$ regularized least squares problem, otherwise known as the Least Absolute Shrinkage and Selection Operator (LASSO). The problem is given by
@@
where $A\in\mathbb{R}^{m\times n},\;u\in\mathbb{R}^{m}$ The corresponding proximal gradient dynamics, referred to as the “firing rate competitive network” [10] are given by
@@
    The LASSO problem is known to satisfy the PL condition [19]. The result follows from Theorems 9. ∎
Example #2: Matrix based problems
Next, consider matrix optimization problems. In this setting, the analogue of sparse vectors are low rank matrices. However, similar to the $\ell_{0}$ norm, imposing a low rank constraint leads to nonconvexity. A convex relaxation for the $\ell_{0}$ norm is the trace norm. The proximal operator of $g(X)=\|X\|_{*}$ is
@@
$\dot{X}=-X+\operatorname{S}_{\alpha\lambda}\bigl{(}{X-\alpha\sum\nolimits_{i=1}^{k}(\operatorname{tr}(A_{i}^{\top}X)-y_{i})A_{i}}\bigr{)},$ (19)
A second popular matrix based optimization problem is the matrix factorization problem, with applications in overparameterized linear neural networks with mean square error loss [25]. In general, this problem is nonconvex. To promote a low-rank solution, it is customary to add terms proportional to the nuclear norm and write the problem as,
@@
Example #3: Sparse neural network training
We conduct a numerical experiment using proximal gradient dynamics to train a feedforward neural network with two hidden layers (dimensions 8 and 2) and a one-dimensional output with $\tanh(\cdot)$ activation on each layer. The network is trained on the scipy moons dataset ($N=200$) for classification, using an $\ell_{1}$ regularizer and binary cross-entropy loss. For the purposes of visualization, we vary two first-layer weights, freezing the other weights, and show the evolution of the dynamics in Figure 1, overlaid on a loss landscape. Since $f$ is differentiable and $g$ is CCP, we observe that the cost is decaying, verifying Theorem 2.333We provide additional implementation details on Github These dynamics could easily be implemented for the purposes of training the whole network. Implementations of such continuous time algorithms on analog circuits could reduce power consumption.
Figure 1: Nonconvex loss landscape of a regularized feed-forward neural network with the trajectory of proximal gradient dynamics. The trajectory goes along a path where the cost function is monotonically decreasing, since the loss function is differentiable and the nonsmooth regularizer is CCP.
@@ VI Discussion
In this letter, we discuss cost based convergence properties of the proximal gradient dynamics. We prove in Theorem 2 that the cost function of the associated optimization problem is nonincreasing for finite-valued regularizers. For non-finite-valued regularizers, we are able to show that the cost is nonincreasing in Theorem 7. Next, we extend the PL conditions to continuous-time proximal gradient dynamics, explicitly showing that the cost function decays exponentially under the PL condition. We use this result to consider the case of parametric optimization problems with time-varying parameters, and provide a cost function tracking bound.
An interesting direction of future work is the application of proximal gradient dynamics to design continuous-time solvers, such as analog circuits and biologically plausible neural networks. Further, these dynamics may also allow a generalized nonsmooth analysis of systems based on energy landscapes. A system that is currently studied under this framework is the human brain [15]. While previous work [24] suggests that the algorithms we design for modeling a brain must be gradient-based, due to our analysis proximal gradient dynamics, one could design biologically plausible algorithms for a nonsmooth energy landscape.
-A Technical Results
Lemma 14 (Proximal optimality condition [21]):
Let $g\colon{}^{n}\rightarrow\overline{\real}$ be CCP. Then for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v}\in{}^{n},\alpha>0$, define $z:=\mathrm{prox}_{\alpha g}(v)$. Then $\alpha^{-1}\bigl{(}v-z\bigr{)}\in\partial g(z).$
Lemma 15 (Forward-invariance of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{Dom}F}$):
  •       Proof:
    {lcss}
    The proof is a consequence of Nagumo’s theorem [7, Theorem 3.1] since $\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))\in\operatorname{Dom}g$ for all $x\in{}^{n}$, and $\operatorname{Dom}g=\operatorname{Dom}F$. {arxiv} First, note that $\operatorname{Dom}F$ is a closed and convex set and that $\operatorname{Dom}F=\operatorname{Dom}g$. Nagumo’s theorem [7, Theorem 3.1], states that the $\operatorname{Dom}g$ is forward invariant provided that $\operatorname{F_{\mathrm{prox}}}(x)$ lies in the tangent cone to $\operatorname{Dom}F$ at $x$ for all $x\in\operatorname{Dom}F$. By definition of $\mathrm{prox}$, we have that $\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))\in\operatorname{Dom}g=\operatorname{Dom}F$ for all $x\in{}^{n}$ and that $z-x$ lies in the tangent cone of $\operatorname{Dom}F$ at $x$ for $z,x\in\operatorname{Dom}F$. Thus, the result is proved. ∎
Lemma 16:
Let $g\colon{}^{n}\rightarrow\real$ be a CCP function and let $\{x_{k}\}_{k=1}^{\infty}\subseteq{}^{n}$ be a sequence such that $x_{k}\to x$ as $k\to\infty$. Then for any $v\in{}^{n}$,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\limsup_{k\to\infty}s_{x_{k}}^{\top}v\leq\max\{s^{\top}v\,|\,s\in\partial g(x)\},}$ (22)
where the sequence $\{s_{x_{k}}\}_{k=1}^{\infty}$ is given by $s_{x_{k}}=\underset{s\in{}^{n}}{\textup{arg\,min}}\|s\|_{2}^{2}$ s.t. $s\in\partial g(x_{k})$.
  •       Proof:
[…]
-B Proof of Theorem 9
Note that if $V(x(0))$ is infinity, there is nothing to show. If $V(x(0))$ is finite, then $V(x(t))$ is finite for all $t\geq 0$ by Lemma 15. Similar to the proof of Theorem 2, consider the Dini derivative of ${V(x(t))=f(x(t))+g(x(t))-F^{*}}$. Using the arguments presented in Theorem 2 upto Equation (11),
$D^{+}V(x)\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$ (23)
where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$. Also, note that
[…]
Therefore,
$\nabla f(x)^{\top}(z-x)+g(z)-g(x)+\frac{1}{2\alpha}\|x-z\|_{2}^{2}$
$\quad=-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha).$
Adding and subtracting $\frac{1}{2\alpha}\|x-z\|_{2}^{2}$ in (23),
$D^{+}V(x)$ $\leq-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha)-\frac{1}{2\alpha}\|x-z\|_{2}^{2}.$ (24)
Using the proximal PL condition (7),
$D^{+}V(x)$ $\leq-\alpha\mu(F(x)-F(x^{\star}))-\frac{1}{2\alpha}\|x-z\|_{2}^{2}.$ (25)
Now, the second term may be bounded above by 00. Using Grönwall’s lemma we get our required result.